f(x)+3x^2-18x+2
I need to graph this but don't know how
Help! I'm lost

Found 2 solutions by nyc_function, lwsshak3:
Answer by nyc_function(2733)

Let y = f(x)
Now we have y = 3x^2-18x+2
The graph of this function will be an upward parabola.
To graph, select 3 negative values for x and then 3 positive values for x.
You pick the values desired for x.
Anyway, here it is:

Let x = -3, -2, and -1. Then solve for y to find 3 points in the form (x,y).
Then let x = 1, 2, and 3. Afterward, solve for y to find 3 points in the form
(x,y).
All you have to do next is connect all 6 points with a smooth curve.

Understand?




Answer by lwsshak3(6505)

I need to graph this
f(x)=3x^2-18x+2
Since this is a second degree equation, the graph is a parabola.
complete the square
y=3(x^2-6x+9)+2-27
y=3(x-3)^2-25
This is an equation of a parabola of the standard form: y=A(x-h)^2+k, with (h.k) being the (x,y) coordinates of the vertex.
For given parabola:
Vertex:(3,-25)
axis of symmetry: x=3
Parabola opens upwards (lead coefficient>0)
y-intercept
set x=0
f(0)=3(3)^2-25=27-25=2
This gives you points at (0,2) and (6,2) because of axis of symmetry.
You now have all the information you need to graph the given function